Amit does a work in 90days ,bittu in 40days and chintu in 12days .they work one after another for a day each, starting with Amit followed by bittu and then by chintu. If the total wages recently after the completion of the work are rs 360 and Amit bittu and chintu share them in the ratio of work done find the respective individuals wages
Answer by Cromlix(4381)

Hi there,
90:40:12
9:4:1.2
Add up = 14.2
Amit's wages = 9/14.2 x 360 = 228.17
Bittu's wages = 4/14.2 x 360 = 101.41
Chintu's wages = 1.2/14.2 x 360 = 30.42
Hope this helps :-)